OpenStudy (anonymous):

factorise: (x^2 + y^2 + z^2) ^ 2 - 4 x^2 y^2

OpenStudy (anonymous):

This is what i wanted to say:- \[(x ^{2} + y ^{2} + z ^{2})^{2} - 4x ^{2}y ^{2}\]

OpenStudy (anonymous):

\[\left( x ^{2}+y ^{2}+z ^{2} \right)^{2}-\left( 2xy \right)^{2}\] \[use a ^{2}-b ^{2}=\left( a+b \right)\left( a-b \right)\]

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

if you want complex factors then it can be factorised further.

OpenStudy (anonymous):

i have to more questions...would u like to help me out?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[4^{x} - 3^{x-1/2} = 3^{x+1/2} - 2^{2x-1}\] solve for x

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

sorry openstudy crashed

OpenStudy (anonymous):

\[\left( 2^{2} \right)^{x}-3^{x}*3^{\frac{ -1 }{2 }}=3^{x}*3^{\frac{ 1 }{2 }}-2^{2x}*2^{-1}\]

OpenStudy (anonymous):

\[2^{2x}-\frac{ 3^{x} }{ \sqrt{3} }=\sqrt{3}*3^{x}-\frac{ 2^{2x} }{2 }\]

OpenStudy (anonymous):

\[2^{2x}\left( 1+\frac{ 1 }{ 2 } \right)=3^{x}\left( \sqrt{3}+\frac{ 1 }{ \sqrt{3} } \right)\]

OpenStudy (anonymous):

\[\frac{ 3 }{ 2 }2^{2x}=\frac{ 4 }{\sqrt{3} }3^{x}\]

OpenStudy (anonymous):

\[\frac{ 2^{2x} }{ 3^{x} }=\frac{ 4 }{ \sqrt{3} }*\frac{ 2 }{3 }\]

OpenStudy (anonymous):

\[\left( \frac{ 2 }{ \sqrt{3} } \right)^{2x}=\left( \frac{ 2 }{ \sqrt{3} } \right)^{3}\]

OpenStudy (anonymous):

2x=3 x=3/2

OpenStudy (anonymous):

if you want any clearification ,you can ask

OpenStudy (anonymous):

i've understood...thnk u so much :) last one: if \[\sqrt{a-b} = (b/a )^{1-2x}\] find x